OpenStudy (anonymous):

Please help me with this!!!! I will give you a medal! My first born! Anything! \[\frac{ \sqrt{a} +2\sqrt{y} }{\sqrt{a}-2\sqrt{y} }\]

OpenStudy (lucaz):

\[\frac{\sqrt{a}+2\sqrt{y}}{\sqrt{a}-2\sqrt(y)}\]

OpenStudy (lucaz):